i do not know if this was the right topic for this question but i guessed.
i am sorry but this is a matching question so there are 4 questions in one kind of.
match each statment on the left with a solution on the right.
1. |a-5|less than 3 a. a=2 or a=8
2. |a-5| greater than 3 b. a is less than 2 or a is greater than 8
3. |a-5|equal to 3 c. a is greater than 2 and a=8
4. |a-5|less than -3 d. there is no solution

Found 2 solutions by Earlsdon, stanbon:
Answer by Earlsdon(6103)

1. |a-5| < 3
a-5 < 3 or a-5 > -3
a < 8 or a >2
2. |a-5| > 3
a-5 > 3 or a-5 < -3
a > 8 or a < 2
3. |a-5| = 3
a-5 = 3 or a-5 = -3
a = 8 or a = 2
4. |a-5| < -3
a-5 < -3 or a-5 > 3
a < 2 Not a solution.
a > 8 Not a solution.

Answer by stanbon(48558)

match each statment on the left with a solution on the right.
1. |a-5|less than 3 a. a=2 or a=8
Ans: no solution
---------------------
2. |a-5| greater than 3 b. a is less than 2 or a is greater than 8
Ans: a is less than 2
----------------------
3. |a-5|equal to 3 c. a is greater than 2 and a=8
Ans: a=8
-----------------------
4. |a-5|less than -3 d. there is no solution
Ans: correct, there is no solution
====================
Cheers,
Stan H.